### v1

# Streaming Technologies and Serialization Protocols: Empirical Performance Analysis

Abstract
Efficiently streaming high-volume data is essential for real-time data analytics, visualization, and AI and machine learning model training. Various streaming technologies and serialization protocols have been developed to meet different streaming needs. Together, they perform differently across various tasks and datasets. Therefore, when developing a streaming system, it can be challenging to make an informed decision on the suitable combination, as we encountered when implementing streaming for the UKAEA’s MAST data or SKA’s radio astronomy data. This study addresses this gap by proposing an empirical study of widely used data streaming technologies and serialization protocols. We introduce an extensible and open-source software framework to benchmark their efficiency across various performance metrics. Our findings reveal significant performance differences and trade-offs between these technologies. These insights can help in choosing suitable streaming and serialization solutions for contemporary data challenges. We aim to provide the scientific community and industry professionals with the knowledge to optimize data streaming for better data utilization and real-time analysis.
Index Terms:
Data streaming, messaging systems, serialization protocols, web services, performance evaluation, empirical study, and applications.
I Introduction
With the exponential increase in data generation from large scientific experiments and the concurrent rise of data-intensive machine learning algorithms within scientific computing [1], traditional methods of data transfer are becoming inadequate. The trend necessitates efficient data streaming methods that allow end-users to access subsets of data remotely. Additionally, the drive for FAIR and open data [2, 3] mandates that such data are, ultimately, publicly accessible to end users over a wide-area network connection.
The Mega-Ampere Spherical Tokamak (MAST) [4] was a spherical tokamak in operation at the UK Atomic Energy Authority (UKAEA), Culham Centre for Fusion Energy (CCFE) from 1999 to 2013 and its upgraded successor was the MAST-U [5] began operation in 2020. These facilities generate gigabytes of data per experimental shot, accumulating substantial data daily.
The lack of public accessibility to the historical archive of data produced by the MAST experiment has limited collaborative opportunities with international and industry partners. The pressing need to facilitate real-time data analysis [6] and leverage recent advancements in machine learning [7] further emphasizes the necessity for efficient data streaming technologies. These technologies must not only handle the sheer volume of data but also integrate seamlessly with analytical tools.
In this paper, we extend the work conducted in [8] for the SKA’s radio astronomy data streaming and visualization. We explore an array of streaming technologies available. We consider the combination of two major choices of technology when implementing a streaming service: (a) the choice of a streaming system, which performs the necessary communication between two endpoints, and (b) the choice of encoding used to convert the data into transmittable formats. Our contributions are as follows:
  • We provide a comprehensive review of 11 streaming technologies and 11 encoding methods, categorized by their underlying principles and operational frameworks.
  • We introduce an extensible software framework designed to benchmark the efficiency of various combinations of streaming technology and serialization protocols, assessing them across 11 performance metrics.
  • By testing 132 combinations, we offer a detailed comparative analysis of their performance across six different data types.
  • Our findings not only highlight the performance differentials and trade-offs between these technologies, but we also discuss the limitations of this study and potential directions for further research.
Through this comprehensive study, we aim to equip the scientific community with deeper insights into choosing appropriate streaming technologies and serialization protocols that can meet the demands of modern data challenges.
Section II briefly reviews the related work in this area. Section III provides an overview of the different serialization protocols and data streaming technologies reviewed in this study. Section IV outlines our experimental methodology, including the performance metrics considered, implementation details of our benchmark framework and the choice of datasets used for evaluation. Section V discusses the results of our experiments across all performance metrics and datasets. Finally, sections VI and VII reflect on the results of our study and draw recommendations of technology choices.
II Related Work
Khan et al. [8] evaluated the performance of streaming data and web-based visualization for SKA’s radio astronomy data. They also conducted a limited analysis on the serialization, deserialization, and transmission latency of two protocols - ProtoBuf and JSON. Our work builds on their research by covering a more extensive range of combinations.
Proos et al. [9] consider the performance of three different serialization formats (JSON, Flatbuffers, and Protobuf) and a mixture of three different messaging protocols (AMQP, MQTT, and CoAP). They evaluate the performance using real “CurrentStatus” messages from Scania vehicles as JSON data payload data. They monitor communication between a desktop computer and a Raspberry Pi. They consider numerous evaluation metrics such as latency, message size, and serialization/deserialization speed.
The authors of [10] compare 10 different serialization formats for use with different types of micro-controllers and evaluate the size of the payload from each method. They test performance with two types of messages 1) JSON payloads obtained from “public mqtt.eclipse.org messages” and 2) object classes from smartphone-centric studies [11, 12].
Fu and Zhang [13] presented a detailed review of different messaging systems. They evaluate each method in terms of throughput and latency when sending randomly generated text payloads. They evaluate each method only on the local device to avoid bias from any network specifics. Orthogonal to our work, they are focused on evaluating the scaling of each system over a number of producers, consumers, and message queues.
Churchill et al. [6] explored using ADIOS2 [14] for transferring large amounts of Tokamak diagnostic data from the K-STAR facility in Korea to the NERSC and PPPL facilities in the USA for near-real-time data analysis.
We differentiate our study from these related works by evaluating 1) A wide variety of different streaming technologies, both message broker-based and RPC-based. 2) considering a large number of data serialization formats, including text, binary, and protocol-based formats. 3) We evaluate the combination of these technologies, developing an extensible framework for measuring and comparing serialization and streaming technologies. 4) Evaluating the performance over 10 different metrics. We comprehensively evaluate 10 different streaming technologies with 12 different serialization methods over 8 different datasets.
III Background
In this paper, we study how the choice of streaming technologies and serialization protocols critically affects data transfer speed. Specifically, we analyze the application of popular messaging technologies and serialization protocols across diverse datasets used in machine learning. Before discussing our experimental setup and results, this section provides an overview of message systems and serialization protocols suitable for streaming data.
III-A Serialization Protocols
TABLE I: A comparison of various serialization protocols. Type: describes how the method serializes data, whether in text or binary format, or relying on a common protocol. Human readable: indicates whether the serialization scheme is legible to a human reader. Defined schema: specifies whether producer and consumer share a common knowledge of the data format prior to transmission. Code generated schema: states whether the serialization requires code to be generated from a predefined protocol.
Protocol Type Binary Human Readable Defined Schema Code Generated Schema Based On
  XML [15] Text
JSON [16] Text
YAML [17] Text
BSON [18] Binary JSON
UBSON [19] Binary JSON
CBOR [20] Binary MessagePack
MessagePack [21] Binary JSON
Pickle [22] Binary
ProtoBuf [23] Protocol
Thrift [24] Protocol
Capn’Proto [25] Protocol
Avro [26, 12, 10] Protocol
In this section, we provide a brief overview of three different categories of serialization protocol: text formats, binary formats, and protocol formats.
III-A1 Text Formats
Extensible Markup Language (XML) [15] is a markup language and data format developed by the World Wide Web consortium. It is designed to store and transmit arbitrary data in a simple, human-readable format. XML adds context to data using tags with descriptive attributes for each data item. It has been extended to various derivative formats, such as XHTML and EXI.
JavaScript Object Notation (JSON) [16] is another human-readable data interchange format that represents data as a collection of nested key-value pairs. JSON is commonly used for data exchange protocol in RESTful APIs. Due to the smaller payload size, it is often seen as a lower overhead alternative to XML for data interchange.
YAML Ain’t Markup Language (YAML) [17] is a simple text-based data format often used for configuration files. It is less verbose than XML and supports advanced features such as comments, extensible data types, and internal referencing.
III-A2 Binary Formats
Binary JSON (BSON) [18] is a binary data format based on JSON, developed by MongoDB. Similar to JSON, BSON also represents data structures using key-value pairs. It was initially designed for use with the MongoDB NoSQL database but can be used independently of the system. BSON extends the JSON format with several data types that are not present in JSON, such as a datetime format.
Universal Binary JSON (UBJSON) [19] is another binary extension to the JSON format created by Apache. UBJSON is designed according to the original philosophy of JSON and does not include additional data types, unlike BSON.
Concise Binary Object Representation (CBOR) [20] is also based on the JSON format. The major defining feature of CBOR is its extensibility, allowing the user to define custom tags that add context to complex data beyond the built-in primitives.
MessagePack [21] is a binary serialization format, again based on JSON. It was designed to achieve smaller payload sizes than BSON and supports over 50 programming languages.
Pickle [22] is a binary serialization format built into the Python programming language. It was primarily designed to offer a data interchange format for communicating between different Python instances.
III-A3 Protocol Formats
Protocol Buffers (ProtoBuf) [23] were developed by Google as an efficient data interchange format, particularly optimized for inter-machine communication. Specifically, ProtoBuf is designed to facilitate remote procedural call (RPC) communication through gRPC [27]. Data structures used for communication are defined in .proto files, which are then compiled into generated code for various supported languages. During transmission, these data structures are serialized into a compact binary format that omits names, data types, and other identifiers, making it non-self-descriptive. Upon receipt, the messages are decoded using the shared protocol buffer definitions.
Thrift [24] is another binary data format developed by Apache Software Foundation or Apache, similar in many respects to ProtoBuf. In Thrift, data structures are also defined in a separate file, and these definitions are used to generate corresponding appropriate data structures in various supported languages. Before transmission, data is serialized into a binary format. Thrift is also designed for RPC communication and includes methods for defining services that use Thrift data structures. However, Thrift has a smaller number of supported data types compared to ProtoBuf.
Capn’Proto [25] is a protocol-based binary format that competes with ProtoBuf and Thrift. Capn’Proto differentiates itself with two main features. First, its internal data representation is identical to its encoded representation, which eliminates the need for a serializing step. Second, its RPC service implementation offers a unique feature called “time travel” enabling chained RPCs to be executed as a single request. Additionally, Capn’Proto offers a byte-packing method that reduces payload size, albeit with the expense of some increase in serialization time. In our experiments, we refer to the byte packed version of Capn’Proto as ”capnp-packed” to differentiate it from the unpacked version, ”capnp”.
Avro [26] is a schema-based binary serialization technology developed by Apache. Avro uses JSON to define schema data structures and namespaces. These schemas are shared between both producer and consumer. One of Avro’s key advantages is its dynamic schema definition, which does not require code generation, unlike competitors such as ProtoBuf. Avro messages are also self-describing, meaning they can be decoded without needing access to the original schema.
We also considered the PSON format [28] and Zerializer [29]. PSON is a binary serialization format with a current implementation limited to C++ and lacks Python bindings, which restricts its applicability for our study. Zerializer, on the other hand, necessitates a specific hardware setup for implementation, placing it outside the scope of our study due to practical constraints. Consequently, while these formats might offer potential advantages, their limitations in terms of language support and hardware requirements precluded their inclusion in our experimental evaluation.
A summary of serialization protocols can be found in Table I. The text-based formats represent data during a text-based markup. While human-readable, text-based formats suffer from larger payload and serialization costs due to the overhead of the markup describing the data. In contrast, binary formats serialize the data to bytes before transmission. These formats are not human-readable, but achieve a better payload size with lower serialization costs. Protocol-based formats also encode data in binary, but differ in that they rely on a predefined protocol definition shared between sender and receiver. Using a shared protocol frees more information out of the transmitted packet, yielding smaller payloads and faster serialization time.
III-B Data Streaming Technologies
TABLE II: A comparison of different data streaming technologies.
Name Type Queue Mode Consume Mode Broker Architecture Delivery Guarantee Order Guarantee Code Generated Protocol Multiple Consumer
ActiveMQ [30] Messaging Pub/Sub & P2P Pull controller-worker at-least-once queue-order
Kafka [31] Messaging P2P Pull P2P All partition-order
Pulsar [32] Messaging P2P Push P2P All global-order
RabbitMQ [33] Messaging Pub/Sub Push/Pull controller-worker at-least/most-once None
RocketMQ [34] Messaging Pub/Sub Push/Pull controller-worker at-least-once queue-order
Avro [35] RPC P2P Pull Brokerless None global-order
Capn’Proto [36] RPC P2P Pull Brokerless None global-order
gRPC [37] RPC P2P Pull Brokerless None global-order
Thrift [38] RPC P2P Pull Brokerless None global-order
ZMQ [39] Low Level P2P Pull Brokerless None global-order
ADIOS2 [14] Low Level P2P Pull Brokerless None global-order
In this section, we discuss three different categories of data streaming technologies: message queue-based, RPC-based, and low-level.
III-B1 Message Queues
ActiveMQ [30], developed in Java by Apache, is a flexible messaging system designed to support various communication protocols, including AMQP, STOMP, REST, XMPP, and OpenWire. The system’s architecture is based on a controller-worker model, where the controller broker is synchronized with worker brokers. The system operates in two modes: topic mode and queuing mode. In topic mode, ActiveMQ employs a publish-subscribe (pub/sub) mechanism, where messages are transient, and delivery is not guaranteed. Conversely, in queue mode, ActiveMQ utilizes point-to-point messaging approach, storing messages on disk or in a database to ensure at-least-once delivery. For our experiments, we utilize the STOMP communication protocol.
Kafka [31] is a distributed event processing platform written in Scala and Java; initially developed by LinkedIn and now maintained by Apache. Kafka leverages the concept of topics and partitions to achieve parallelism and reliability. Consumers can subscribe to one more topic, with each topic divided into multiple partitions. Each partition is read by a single consumer, ensuring message order within that partition. For enhanced reliability, topics and partitions are replicated across multiple brokers within a cluster. Kafka employs a peer-to-peer (P2P) architecture to synchronize brokers, with no single broker taking precedence over other brokers. Zookeeper [40] manages brokers within the cluster. Kafka uses TCP for communication between message queues and supports only push-based message delivery to consumers while persisting messages to disk for durability and fault tolerance.
RabbitMQ [33], developed by VMWare, is a widely used messaging system known for its robust support for various messaging protocols, including AMQP, STOMP, and MQTT. Implemented in Erlang programming language, RabbitMQ leverages Erlang’s inherent support for distributed computation, eliminating the need for a separate cluster manager. A RabbitMQ cluster consists of multiple brokers, each hosting an exchange and multiple queues. The exchange is bound to one queue per broker, with queues synchronized across brokers. One queue acts as the controller, while the others function as workers. RabbitMQ supports point-to-point communication and both push and pull consumer modes. Although message ordering is not guaranteed, RabbitMQ provides at-least-once and at-most-once delivery guarantees. RabbitMQ faces poor scalability issues due to the need to replicate each queue on every broker. Our experiments utilize the STOMP protocol for communication with the pika python package.
RocketMQ [34], developed by Alibaba and written in Java, is a messaging system that employs a bespoke communication protocol. It defines a set of topics, each internally split into a set of queues. Each queue is hosted on a separate broker within the cluster, and queues are replicated using a controller-worker paradigm. Brokers can dynamically register with a name server, which manages cluster and query routing. RocketMQ guarantees message ordering, and supports at-least-once delivery. Consumers may receive messages from RocketMQ either using push or pull modes. Message queuing is implemented using the pub/sub paradigm, and RocketMQ scales well with a large number of topics and consumers.
Pulsar [32], created by Yahoo and now maintained by Apache, is implemented in Java and designed to support a large number of consumers and topics while ensuring high reliability. Pulsar’s innovative architecture separates message storage from the message broker. A cluster of brokers is managed by a load balancer (Zookeeper). Similar to Kafka, each topic is split into partitions. However, instead of storing messages within partitions on the broker, Pulser stores partition references in bookies. These bookies are coordinated by a bookkeeper, which is also load-balanced using Zookeeper. Each partition is further split into several segments and distributed across different bookies. The separation of message storage from message brokers means that if an individual broker fails, it can be replaced with another broker without loss of information. Similarly, if a bookie fails, the replica information stored in other bookies can take over, ensuring data integrity. Pulsar’s architecture allows it to offer a global ordering and delivery guarantee, although this high reliability and scalability come at the cost of extra communication overhead between brokers and bookies.
For a detailed overview of different message queue technologies, please refer  [13].
III-B2 RPC Based
gRPC [37], developed by Google, is an RPC framework that utilizes ProtoBuf as its default serialization protocol. To define the available RPC calls for a client, gRPC requires a protocol definition written in ProtoBuf. While ProtoBuf is the standard, sending arbitrary bytes from other serialization protocols over gRPC is possible by defining a message type with a bytes field. The Python gRPC implementation supports synchronous and asynchronous (asyncio) communication. For all our experiments with gRPC, we use asynchronous communication.
Capn’Proto [36] and Thrift also have their own RPC frameworks. Similar to gRPC, these frameworks define remote procedural calls within their protocol definitions, using their own syntax specification. Like gRPC, they allow the transmission of arbitrary bytes by defining a message with a bytes field.
Avro provides RPC-based communication protocol as well. Unlike other RPC-based methods, Avro does not require the RPC protocol to be explicitly defined. This flexibility comes at the expense of stricter type validation, setting Avro apart from systems such as gRPC and Thrift.
III-B3 Low Level
In addition to RPC and messaging systems, we consider two low-level communication systems: ZeroMQ and ADIOS2. Like RPC systems, they do not rely on an intermediate broker for message transmission.
ZeroMQ (ZMQ) [39] is a brokerless communications library developed by iMatix. It is a highly flexible message framework that uses TCP sockets and supports various messaging patterns, such as push/pull, pub/sub, request/reply, and many more. Notably, ZeroMQ’s zero-copy feature minimizes the copying of bytes during data transmission, making it well-suited for handling large messages. In our experiments, we implement a simple push/pull messaging pattern to avoid the additional communication overhead associated with RPC methods.
The ADaptable Input Output System (ADIOS) [14] is a unified communications library developed as part of the U.S Department of Energy’s (DoE) Exascale Computing Project. It is designed to stream exascale data loads for interprocess and wide area network (WAN) communication. In this study, we compare the WAN capabilities of ADIOS, which uses ZeroMQ for it’s messaging protocol. We use ADIOS2 for communication and the low-level Python API to facilitate communication between client and server.
We do not consider other RPC systems such as Apache Flight. Instead, we rely on ProtoBuf and gRPC for their communication protocols.
A summary of the comparison of various data streaming technologies can be found in Table II. Message queue-based technologies use message queues and a publish/subscribe model to transmit data. Producers publish messages to a topic, and multiple consumers can subscribe to these topics to read messages from the queue. These systems operate in push mode, where the system delivers messages to consumers, or in pull mode, where consumers request messages from the message queuing system. RPC-based technologies define a communication protocol shared between producers and consumers, eliminating the need for an intermediate broker. Producers respond to remote procedure calls (consumer requests) to provide data. Low-level communication protocols and the ADIOS also do not require an intermediate broker. Unlike RPC technologies, they do not wait for clients’ requests to send messages, reducing communication overhead. ZeroMQ and ADIOS support zero-copy messaging transfer of raw bytes, which is particularly beneficial for large array workloads where encoding and copying data can be costly.
These technologies differ in their fault tolerance. Message queuing systems prioritize reliability by caching messages to disk to prevent load shedding during high message rates. In contrast, RPC systems keep all requests in memory, offering faster performance at the expense of lower fault tolerance. Many protocol-based serialization formats introduced earlier include RPC communications libraries that support sending arbitrary bytes. For example, Protobuf-encoded messages can be sent using Avro RPC communication library.
IV Empirical Study Design
Figure 1: Illustrates the data flow from producer to consumer, indicating the places at which various performance metrics are recorded. These metrics include (A) $L_{o}$: object creation latency, (B) $T_{o}$: object creation throughput, (C) $C$: compression ratio, (D) $L_{s}$: serialization latency, (E) $L_{d}$ deserialization latency, (F) $T_{s}$: serialization throughput, (G) $T_{d}$: deserialization throughput, (H) $L_{trans}$: transmission latency, (I) $T_{trans}$: transmission throughput, (J) $L_{tot}$: total latency, and (K) $T_{tot}$: total throughput.
Figure 2: Diagram showing the architecture of our streaming framework. A Runner is used to create a Producer and Consumer pair for each type of streaming technology. Both producer and consumer are instantiated with a Marshaler that encodes data to the desired format (e.g. JSON, ProtoBuf etc.). Producers are created with a data stream object that generates data samples for transmission. Depending on the streaming method, the Consumer and Producer may connect to an external message broker.
The objective of this empirical study is to investigate and compare various streaming technologies and serialization protocols for scientific data. We examine the interplay between serialization protocol and streaming technology by exploring different combinations of them. We conduct experiments on all the technologies discussed in section III, which includes 11 different streaming technologies and 15 different serialization protocols. We test each combination of technology across eight different payloads, resulting in $11\times 15\times 8=1320$ different combinations.
IV-A Performance Metrics
We consider 11 performance metrics: seven of these metrics are associated with each serialization protocol, and the remaining four are linked to the combination of streaming technology and serialization protocol. To define the metrics, we first need to establish the different sizes of data as it passes through our pipeline. We denote the size of the data straight from the stream as $S_{d}$, the size of the data after object creation as $S_{o}$, and the size of the payload after encoding to bytes as $S_{p}$. Additionally, we define the number of samples in a dataset as $N$.
To evaluate the performance of each serialization technology we measure:
  1. 1.
    Object Creation Latency ($L_{o}$) – This measures the total time taken to convert the program-specific native format (e.g. numpy array, xarray dataset) into the format required for transmission. This is an important metric because some formats, such as Capn’Proto store their data internally in a serialization-ready format. However, in reality, we often need to work with arrays that are in an analysis-ready format, such as numpy array or xarray dataset. Converting between the two models naturally incurs a penalty since it involves copying the data.
  2. 2.
    Object Creation Throughput ($T_{o}=\frac{S_{d}N}{\sum L_{o}^{(i)}}$) – This is similar to serialization and deserialization throughput, we consider object creation throughput is a sum of latencies for all samples sent) of converting a native object (e.g., numpy array, xarray dataset) to the transmission format expected by the protocol (e.g., a ProtoBuf object or a Capn’Proto object).
  3. 3.
    Compression Ratio ($C=\frac{S_{p}}{S_{o}}\times 100$) – This is defined as the ratio of the size of the payload $S_{p}$ after serialization to the size of the object $S_{o}$, for a given encoding. A smaller compression ratio ultimately means less data to be transmitted over the wire, and therefore, protocols that produce a smaller payload should be more performant.
  4. 4.
    Serialization Latency ($L_{s}$) – This is the total time taken in seconds to encode the original data into the serialized format for transmission. Encoding data with any serialization protocol incurs a non-zero cost due to the need to format, copy, and compress data for transmission. A larger serialization time can potentially negate the benefits of a smaller payload size because it increases the total transmission time.
  5. 5.
    Deserialization Latency ($L_{d}$) – This is similar to serialization time, this metric measures the total time required to deserialize a payload after transmission across the wire. As with serialization time, a slow deserialization time can also negate the effects of a smaller payload.
  6. 6.
    Serialization Throughput ($T_{s}=\frac{S_{o}}{L_{s}}$) – This is the serialization time divided by the size of the object to be transmitted. This measures how many bytes per second a serialization protocol can handle, independent of the size of the data stream.
  7. 7.
    Deserialization Throughput ($T_{d}=\frac{S_{o}N}{\sum L_{d}^{(i)}}$) – This is the deserialization time divided by the size of the object received. This measures how many bytes per second a deserialization protocol can handle, independent of the size of the data stream.
For streaming technologies, we consider two different performance metrics:
  1. 8.
    Transmission Latency ($L_{trans}$) – This is the time taken for a payload to be sent over the wire, excluding the time taken to encode the message.
  2. 9.
    Transmission Throughput ($T_{trans}=\frac{S_{d}N}{\sum L_{trans}^{(i)}}$) – This is similar to total throughput, but considers the payload size divided by the time taken to send the message over the wire, exclusive of the serialization time.
  3. 10.
    Total Latency ($L_{tot}$) – This is the total time for a payload to be transmitted from producer to consumer, inclusive of the serialization time.
  4. 11.
    Total Throughput ($T_{tot}=\frac{S_{d}N}{\sum L_{tot}^{(i)}}$) – This is the original data object size divided by the total time to send the message. Throughput measures the rate of bytes that can be communicated over the wire.
Finally, we also investigate the effect of batch size on the throughput. Grouping data into batches is a common requirement during machine learning training, and we show increasing the batch size while lowering the number of communications has a positive effect on throughput.
We make a distinction between transmission time and total time (Fig. 1). The total time is the end-to-end transmission of a message, including the time to serialize the message and send it over the wire. Transmission time is the time taken to transmit the payload excluding the serialization and deserialization times. Similarly, we can calculate total and transmission throughput.
IV-B Dataset
In our experiments, we consider eight different payloads, ranging from simple data to common machine learning workloads, and include fusion science data. Our goal is to cover a range of scenarios. This section briefly describes the datasets used to evaluate performance with various streaming technologies and serialization protocols.
  • Numerical Primitives: As a baseline comparison, we use simple datasets consisting of randomly generated numerical primitives for int32, float32, and float64 types.
  • BatchMatrix: A synthetic dataset where each message consists of a randomly generated 3D tensor of type float32 with shape $\{32,100,100\}$ to simulate sending a batched set of image samples.
  • Iris Data: This is a dataset using the well-known Iris dataset [41]. The Iris dataset contains an array of four float32 features and a one-dimensional string target variable.
  • MNIST: We use the widely used MNIST machine learning image dataset [42] as a realistic example of streaming 2D tensor data.
  • Scientific Papers: The scientific papers dataset is a well-known dataset in the field of NLP and text processing [43]. The dataset comprises 349,128 articles of text from PubMed and arXiv publications. Each sample is repeated as a collection of string for properties such as article, abstract, and section names for transmission.
  • Plasma Current Data: A more realistic example of scientific data, we use plasma current data from the MAST tokamak [4]. Each set of plasma current data contains three 1D arrays of type float32: data, time, and errors. The “data” array represents the amount of current at each timestep, the “time” represents the time the measurement was taken in seconds, and the “errors” represents the standard deviation of the error in the measured current.
IV-C Implementation and Experimental Setup
We developed a framework to measure the performance of streaming and serialization technology. The architecture diagram of our framework is shown in Figure 2, which follows service-oriented architecture [44, 45] and is implemented in Python. We used the appropriate Python client library for each streaming and serialization technology. The source code can be found in our GitHub repository [46].
The user interacts with the framework through a command-line interface. A test runner sets up both the server-side and client-side of the streaming test.
The server side requires the configuration of three components:
  • DataStream: component handles loading data for transmission. This can be any one of the payloads described in section IV-B.
  • Producer: functions as the server side of the application. It packages data from the selected data stream and transmits it over the wire using the selected streaming technology, which may be any of the technologies described in section III-B.
  • Marshaler: handles the serialization of the data from the stream using the specified serialization protocol. This can be any of those described in section III-A.
The configuration of the client side is similar but only requires a marshaler to be configured to match the one used for the producer. It does not require knowledge of the data stream.
  • Consumer: functions as the client side of the application. It receives data transmitted by the producer using the selected streaming technology, processes the incoming messages, and performs the necessary actions. Producers and consumers interact using a configured protocol.
  • Broker: required by the streaming protocol (e.g., for Kafka, RabbitMQ, etc.) are run externally from the test in the background. In our framework, we configure all brokers using docker-compose [47] to ensure that our broker configurations are reproducible for every test.
  • Logger: is used by the marshaler to capture performance metrics for each test in a JSON file. For each message sent, the logger captures four timestamps: 1) before serialization, 2) after serialization, 3) after transmission, and 4) after deserialization. Using these four timestamps, we can calculate the serialization, deserialization, transmission, and total time. Additionally, the logger captures the payload size of each message immediately after serialization. With this additional information, we can calculate the average payload size and throughput of the streaming service.
ADIOS and ZeroMQ can directly send array data without copying the input array. However, to achieve this, the array data must be directly passed to the communication library without serialization. Therefore, we additionally consider ZeroMQ and ADIOS to have their own ”native” encoding strategy for each stream, which is only used with their respective streaming protocol. This allows for a fair comparison with other technologies because sending an encoding array with ADIOS or ZeroMQ incurs an additional copy that could be circumvented by properly using their zero-copy functionality.
Following the convention of previous work [13, 8], we run each streaming test locally, with the producer and consumer on the same machine to avoid network-specific issues.
V Results
In this section we present the results of our experiments with the combination of different streaming technologies, serialization protocols, and data streams.
1) Object Creation Latency – We use different datasets that originate from various data analysis types, like NumPy or Xarray dataset. Depending on the encoding protocol, we may need to copy the data from its native format to a specific format like Capn’Proto or Protobuf objects. This copying process adds some overhead that should be taken into consideration. However, for encoding protocols like JSON, BSON, and Pickle that do not require format changes, we store the data in a Pydantic class. The results in Figure 3 show that for larger array datasets like BatchMatrix, Plasma, and MNIST encoding methods such as Protobuf, Thrift, and Captn’Proto tend to have higher object creation latency as they need to copy data into their own data types.
Figure 3: Object latency ($L_{o}$) measured in milliseconds, for various data streams (x-axis) and encoding methods (color).
2) Object Creation Throughput – We consider the object creation throughput for each serialization method. The object creation time measures the time to convert data from the native data structure (such as a NumPy array) to the serialization format. Object creation time is important to consider if the format that the data will be used in will be different from the format it is sent over the wire. Typically, object creation force a copy of the data to be sent, which impacts the total throughput, especially when considering large array like data.
Figure 4 shows the object creation throughput for each dataset and each encoding method. It is interesting to note here that protocol based methods incur a greater penalty for the object creation. This effect is more noticeable in larger datasets such as the BatchMatrix and Plasma datasets.
Figure 4: Object creation throughput ($T_{o}$) measured in megabytes per second, for various data streams (x-axis) and encoding methods (color).
Figure 5: The compression ratio ($C$), for various data streams (x-axis) and serialization protocols (color).
3) Compression Ratio – The payload size, a crucial performance metric of serialization protocols, is independent of the choice of streaming protocols. Therefore, we have calculated the average compression ratio over all runs for each serialization protocol. Figure 5 presents the results for each protocol and data stream. Notably, Pickle, Avro, and XML consistently produce the largest payload sizes, often exceeding the original size. This is due to the inefficiency of their text-based encodings and the additional meta-data tags they add as overhead. Pickle, a binary format for storing Python objects, is particularly known for its large sizes and is not optimal for encoding data for streaming.
The results show that the serialization protocol, Capn‘Proto, outperforms others in terms of payload size. The packed option of Capn‘Proto, also known as capnp-packed, is responsible for additional size efficiency. The capnp-packed format is closely followed by several binary serialization formats that show similar performance. The reason behind this performance can be attributed to their ability to achieve near-identical compression, which is close to the limits of what is possible for that particular data stream.
Examining across data streams, it can be seen that the BatchMatrix dataset is fundamentally limited. This is because it is made up of randomly generated numbers, making it incompressible due to the lack of redundancy in the data. Conversely, for more realistic data such as MNIST and Plasma, a much higher compression ratio is achieved.. Better compression is achieved for formats such as Capn’Proto Packed, which exploit the redundancy in the data to achieve greater compression. Text-based formats, such as YAML, JSON, XML, and Avro, achieve significantly worse compression in comparison. In fact, due to the extra markup required for these formats, they can produce a larger payload size that the original data.
Figure 6: Serialization ($L_{s}$) latency for different data streams (x-axis) and encoding protocols (color). Protocol encoding methods such as Protobuf and Captn’Proto consistently offer the best performance in terms of both serialization. Text-based encodings (YAML, XML, etc.) add a large latency penalty to serialization by increasing the verbosity of the data.
Figure 7: Deserialization ($L_{d}$) latency for different data streams (x-axis) and encoding protocols (color). As with serialization, protocol encoding methods such as Protobuf and Captn’Proto offer the best performance. Likewise, Text-based encodings (YAML, XML, etc.) add a large latency penalty to serialization by increasing the verbosity of the data.
4) Serialization Latency – The results for serialization time are shown in figure 6. There is a clear trend across all data streams from text based protocols being the slowest (Avro, YAML, etc.) towards binary encoded protocol-based methods (Capn’Proto, protobuf, etc.) being the fastest. Binary encoded but no-protocol methods fall in between these two extremes.
It is interesting to note that Capn’Proto has the fastest serialization time. This is likely due to the fact that Capn’Proto stores data in format that is ready for serialization over the wire.
5) Deserialization Latency – The results for serialization time are shown in figure 7. Again, a clear trend may be seen across all data streams from text based protocols to binary encoded protocol based methods.
Like serialization, Capn’Proto is generally the fastest deserialization methods across all tests. As mentioned above, this is likely due to Capn’Proto storing the data in a pre-serialized form.
6) Serialization Throughput – Figure 8 shows the average throughput for serialization of the data using different types of protocols. It is evident from the graph that serialization techniques based on protocols such as ProtoBuf, Thrift, and Capn’Proto offer the highest serialization throughput. Binary methods that are protocol independent offer moderate throughput performance with the added advantage of greater flexibility as compared to protocol methods. Text-based methods perform the worst due to their high serialization overhead.
Surprisingly, Avro also performs quite well by this metric. We believe that despite being a human-readable text-based method, it is also a protocol-based method. This means that both the producer and consumer are aware of the types and data structures being transmitted over the wire, facilitating faster throughput.
Figure 8: Serialization ($T_{s}$) and deserialization ($T_{d}$) throughput for each encoding averaged over all data streams.
7) Deserialization Throughput – Figure 8 also shows the average throughput for deserialization of the data using different types of protocols. It is noticeable that deserialization throughput time across all methods is smaller, indicating that deserialization is a main bottleneck to transmission.
8) Transmission Latency – Figure 9 shows the transmission latency for various combinations of serialization and streaming technologies. The heatmap in each combination is sorted by the average latency from lowest to highest for each streaming technology.
Across all technologies, it is observed that transmission latency is largely dependent on the choice of streaming technology rather than the choice of serialization protocol. In streaming technologies, a broker is required as an intermediary, which increases the overall latency, whereas RPC technologies have no broker, and hence, have lower latency. Among messaging technologies, RabbitMQ performs better with larger payloads, while ActiveMQ achieves lower latency with smaller payloads but performs worst on the largest payload (e.g., BatchMatrix). In RPC-based methods, Thrift consistently has the lowest latency except for the BatchMatrix stream, where Capn’Proto narrowly beats Thrift.
With larger payloads such as BatchMatrix and plasma data streams, the impact of serialization protocol becomes more noticeable. It is challenging to identify a trend between encoding protocols in terms of latency, except that it is crucial to note the inefficiency of using XML and YAML for larger payloads.
For the BatchMatrix data stream, an issue arises when sending a large YAML-encoded payload through the Python API, which causes ADIOS to produce a segmentation fault. Therefore, subsequent latency and throughput plots result in NaN, the empty cells in Figure 9.
Figure 9: Each subplot or heatmap shows transmission latency ($L_{trans}$) for different serialization protocols (vertical axis) and streaming technologies (horizontal axis). Dark red indicates higher latency.
9) Transmission Throughput – By examining the throughput, we gain better understanding of how different protocols affect transmission. Figure 10 shows that RPC methods achieve higher transmission throughput than message streaming technologies. When dealing with larger payloads, such as the BatchMatrix and Plasma data streams, protocol-based serialization choices such as Thirft, Capn’Proto, and Protobuf provide higher throughput than other methods. Interestingly, messagepack also performs well with larger payloads.
Similar to latency, the choice of streaming technology is more important than the encoding. However, a trend towards protocol encoding methods can be observed on some larger datasets, such as the plasma dataset.
Figure 10: Each subplot or heatmap shows transmission throughput ($T_{trans}$) for different serialization protocols (vertical axis) and streaming technologies (horizontal axis).
10) Total Latency – Figure 11 shows the total latency for all combinations tat were tested. As before, it is clear that the Thrift, Capn’Proto, and ZeroMQ all perform well in these tests. ZeroMQ offers the lower latency in the BatchMatrix test because it avoids the overhead of copying the data into a new structure, as is the case with Thift or Protobuf. Among the broker based methods, RabbitMQ consistently performs well.
When it comes to encoding methods, protocol-based methods generally perform the best across all datasets and streaming methods. However, it is not clear which method offers the lowest latency in general. Protocol-based methods can achieve high throughput by mixing encoding protocols and RPC frameworks. For example, considering the MNIST dataset, Capn’Proto achieves the lowest latency with the thrift protocol.
There is a clear trend towards protocol encoding for complex data sets such as Iris, MNIST, and Plasma. Among streaming technologies, Thrift generally shows the best performance.
Figure 11: Each subplot or heatmap shows total latency ($L_{tot}$) for different serialization protocols (vertical axis), and streaming technologies (horizontal axis).
11) Total Throughput – Figure 12 shows total throughput, which is consistent with the total latency results discussed in the previous section. Protocol-based methods achieve the highest throughput. Among all the serialization protocols, Thrift is generally the best performing one. ZeroMQ performs well with the biggest dataset, BatchMatrix. Although the best encoding method is inconclusive, it shows a trend toward protocol-based methods, which give the highest throughput.
Figure 12: Each subplot or heatmap shows total throughput ($T_{tot}$) for different serialization protocols (vertical axis), and streaming technologies (horizontal axis).
Figure 13: Total throughput $T_{tot}$ for various batch sizes (e.g., ranging from 1 to 256) for the MNIST dataset.
12) Effect of Batch Size on Throughput – In machine learning applications, data is often processed in batches. Our findings underscore the potential of batching data before transmission to enhance throughput. However, it is crucial to note that the batch size can significantly impact the throughput of the method used.
Figure 13 shows the throughput of the MNIST dataset with a variable batch size. As the batch size increases, the throughput reduces. This is due to the increased overhead of copying and serializing data for transmission. However, when the batch size is increased beyond 32 images per batch, the overall throughput begins to improve because fewer packets are needed to be communicated over the network. For binary and protocol encoding methods, increasing the batch size is shown to also increase the throughput. This observation is consistent with previous results; generally, protocol-based methods offer the best throughput. At larger batch sizes ($>128$), the throughput continues to increase because fewer transmission time is significantly slower than the serialization/deserialization cost. So grouping many examples into a single transmission improves throughput.
VI Discussion
We can draw several conclusions based on the experiments presented in this work. We identify the following key points from our results:
VI-A Recommendations
RPC systems are faster than messaging broker systems due to the overhead of the intermediate broker. This makes RPC highly efficient for high throughput and low latency transmission of large data, although they do not offer the same delivery guarantees as message broker systems.
We found that the choice of messaging technology has a greater impact than the encoding protocol.
Protocol-based encoding methods such as Capt’n Proto and ProtoBuf perform best for complex data that can be compressed, while MessagePack is a competitive choice for smaller or random data. Protocol-based encoding methods offer the fastest serialization and best compression, with Thrift offering the best throughput and Capn’Proto offering the best compression. Binary encoding methods offer more flexibility at the cost of slower encoding speed. Among the binary encoding methods we tested, MessagePack generally performed the best. Considering text-based protocols, JSON offered the best performance due to its lightweight markup and smaller payload size compared to YAML or Avro.
Among different messaging technologies, we generally found that Apache Thrift achieves very high throughput and low latency across various scenarios. With message broker systems, RabbitMQ generally demonstrates the best performance.
Surprisingly, we did not observe much of a difference when combining different protocol-based encoding and messaging systems. We hypothesized that ProtoBuf would be most efficient when combined with the gRPC, or that Capn’Proto’s RPC implementation would perform best with Capn’Proto. However, this appears not to be the case.
Larger batch sizes facilitate higher throughput for array datasets, as shown by our throughput and batch size experiment in Figure 13, when using either a binary or protocol-based serialization method. For text-based encoding methods, adding the required markup and lack of compression destroys any advantage of batching data for transmission.
VI-B Limitations and future directions
One notable limitation is that in this study we did not investigate the potential of scaling with multiple clients. Previous research has examined this aspect for message queuing systems [13]. A future study could focus on examining the reliability of various RPC technologies based on the number of consumers.
VII Conclusion
In this work, we investigated 132 combinations of different encoding methods and messaging technologies. We evaluated their performance across 11 different metrics and benchmarked each combination against 6 different datasets, ranging from toy datasets to machine learning, to scientific data from the fusion energy domain. We found that messaging technology has the biggest impact on performance, regardless of over the specific serialization method used. Protocol-based methods offer the highest throughput and lowest latency but at the expense of flexibility and robustness. Protocol encoding methods offered the best performance, at the cost of flexibility. Notably, we did not see much difference when combining different protocol-based encoding and messaging systems. Finally, we found that the batch size affects the data throughput for all binary and protocol-based encoding methods.
Contribution
SJ: Designed and implemented the experimental framework, shaping the research methodology and contributions to the writing and conceptualization of the paper. NC: Provided the MAST data for the study and offered expertise in the fusion domain, enhancing the scientific rigor of this empirical study and editing and refining the manuscript. SK: Provided technical supervision and introduced the core idea, building upon SK’s prior work at the University of Oxford, and contributed to the writing and editing of the paper, figures, and plots.
Acknowledgment
We would like to thank our colleagues at UKAEA and STFC for supporting the FAIR-MAST project. Additionally we would like to thank Stephen Dixon, Jonathan Hollocombe, Adam Parker, Lucy Kogan, and Jimmy Measures from UKAEA for assisting our understanding of the Fusion data. We would also like to extend our thanks to the wider FAIR-MAST project which include Shaun De Witt, James Hodson, Stanislas Pamela, Rob Akers from UKAEA and Jeyan Thiyagalingam from STFC. We also want to extend our gratitude to the MAST Team for their efforts in collecting and curating the raw diagnostic source data during the operation of the MAST experiment.